OpenStudy (anonymous):

If $ 2500 is invested in an account that pays interest compounded continuously, how long will it take to grow to $ 5000 at 3%? Round to the nearest tenth

OpenStudy (anonymous):

If $ 2500 is invested in an account that pays interest compounded continuously, how long will it take to grow to $ 5000 at 3%? Round to the nearest tenth. ** Compound interest formula for interest compounded continuously: A=Pe^rt, P=initial investment, r=annual interest rate, A=amount after t years For given problem: 5000=2500e^.03t e^.03t=5000/2500=2 take log of both sides .03t lne=ln2 .03t*1=ln2 t=ln2/.03=23.1 years is this right